### v2

# Lower-mass-gap Black Holes in Dense Star Clusters

Abstract
The existence of compact stellar remnants in the mass range has long been debated. This so-called lower mass gap was initially suggested by the lack of low-mass X-ray binary observations with accretors about , but it has recently been called into question following newer observations, including a lower-mass-gap candidate with a millisecond pulsar companion in the dense globular cluster NGCĀ 1851. Here we model NGCĀ 1851 with a grid of similar dense star clusters utilizing the state-of-the-art Monte Carlo -body code CMC, and we specifically study the formation of lower-mass-gap black holes. We demonstrate that both massive star evolution and dynamical interactions can contribute to forming lower-mass-gap black holes. In general, the collapse of massive remnants formed through mergers of neutron stars or massive white dwarfs produces the largest number of lower-mass-gap black holes among all formation channels. However, in more massive clusters, supernova core collapse can contribute comparable numbers. Our NGCĀ 1851-like models can reproduce millisecond pulsar ā lower-mass-gap black hole binaries similar to the observed system. Additionally, the lower-mass-gap black holes can also become components of dynamically assembled binaries, and some will be in merging black holeāneutron star systems similar to the recently detected gravitational wave source GW230529. However, the corresponding merger rate is probably .
1 Introduction
The maximum mass of neutron stars (NSs) and the transitional mass between NSs and black holes (BHs) have been long-standing questions in astrophysics. Theoretical calculations indicate that NSs have gravitational masses (e.g., Rhoades & Ruffini, 1974; Kalogera & Baym, 1996), while pulsar timing observations have identified (Cromartie etĀ al., 2020; Fonseca etĀ al., 2021) as the most massive NS to date. At the same time, dynamical mass measurements of BHs in X-ray binaries have suggested a lack of BHs with mass , indicating the possible existence of a mass gap, usually referred to as the ālower mass gapā (henceforth LMG), between and (Bailyn etĀ al., 1998; Ćzel etĀ al., 2010; Farr etĀ al., 2011). These findings inspired searches for the origins of the LMG, including those resulting from core-collapse supernovae (Fryer etĀ al., 2012; Belczynski etĀ al., 2012; Kochanek, 2014; Liu etĀ al., 2021; Fryer etĀ al., 2022) and binary evolution of massive stars and compact objects (e.g., Gao etĀ al., 2022; Siegel etĀ al., 2023).
Recently, however, various observations have revealed several candidates residing in the purported LMG. These include compact objects in non-interacting binaries (Thompson etĀ al., 2019; Jayasinghe etĀ al., 2021), an isolated compact object detected by microlensing (Lam etĀ al., 2022), as well as gravitational wave (GW) sources (GW230529 The LIGO Scientific Collaboration etĀ al., 2024, GW200115 Abbott etĀ al., 2021a, and GW190814 Abbott etĀ al., 2020). These observations suggest that the LMG may not be empty.111However, previous analyses of the mass spectra of GW detections still found evidence for a paucity of objects in the LMG, without conclusively resolving it (e.g., Abbott etĀ al., 2021b; Farah etĀ al., 2022).
Recently, yet another LMG candidate was identified in the Galactic globular cluster NGCĀ 1851 (Barr etĀ al., 2024). This object, with inferred mass (95% confidence interval), was found via radio timing of its companion, the millisecond pulsar PSR J0514ā4002E (NGCĀ 1851E). This could be the most massive NS or least massive BH detected in a millisecond pulsar (MSP) binary to date. MSPs are abundant in globular clusters (GCs) and are understood to form at high rates in these systems due to the role of dynamical processes enabled by high stellar densities. As pointed out by Barr etĀ al. (2024), the companion of NGCĀ 1851E bears a striking resemblance in mass to the remnant of the known double neutron star binary merger, GW170817, observed by LIGO-Virgo-KAGRA (LVK; Abbott etĀ al., 2017), potentially hinting that this object formed via a similar mechanism. The double NS merger origin is also motivated by the presence of a known double NS binary with an inspiral time of roughly Myr in another Galactic globular cluster, M15 (Jacoby etĀ al., 2006). Additional formation scenarios are also viable including mergers of massive white dwarf binaries and accretion-induced collapse via binary Roche lobe overflow or tidal disruption of stars (e.g., Kremer etĀ al., 2021, 2022; Ye etĀ al., 2024). Furthermore, scenarios involving single-star evolution cannot be ruled out due to the uncertainties in supernova natal kicks (e.g., OāDoherty etĀ al., 2023, and references therein). However, retention within a GC with central escape velocity places constraints on the natal kick such an object could have received at birth (e.g., Ye etĀ al., 2019, and references therein).
It is notable that the host cluster NGCĀ 1851 has one of the highest central densities of Milky Way GCs (e.g., Baumgardt & Hilker, 2018). Indeed, previous studies of the dynamics of NSs in GCs have demonstrated that clusters with high central densities (especially those that have undergone cluster core collapse) feature a markedly enhanced rate in NS dynamics (e.g., Ye etĀ al., 2019). In addition to the NGCĀ 1851E binary, clusters that have undergone (or are near to) core collapse are hosts to the aforementioned double NS binary M15C, a population of six apparently young radio pulsars (Boyles etĀ al., 2011; Kremer etĀ al., 2023), and an overabundance of single MSPs without binary companions (Ye etĀ al., 2024), all pointing to processes uniquely enabled by cluster dynamics.
In this study, we explore the formation and dynamics of LMG BHs in dense stellar clusters, similar to the companion of NGCĀ 1851E. We focus in particular upon the formation of these objects via NS collapse, triggered by several mechanisms enabled by dynamical interactions including double NS mergers, white dwarf mergers, and stellar tidal disruptions. We then explore the subsequent dynamical evolution of these mass-gap objects in clusters including their interactions with BHs, NSs, and other stars. We demonstrate that objects like the NGCĀ 1851E binary can form naturally in clusters like NGCĀ 1851 and predict other related GW/electromagnetic sources that may potentially be observed in the future.
We describe the Monte Carlo -body code used for simulating GCs and the compact objects within in SectionĀ 2. We present simulations matching the observed properties of NGCĀ 1851 and describe the grid of core-collapsed cluster simulations in SectionĀ 3. In SectionĀ 4, we explore the formation of LMG BHs through massive star evolution and close dynamical encounters in dense star clusters. We study the formation of LMG binaries in SectionĀ 5, and potential GW sources and detectable transient signals consisting of LMG BHs in SectionĀ 6. Lastly, we discuss the model uncertainties in SectionĀ 7 and conclude in SectionĀ 8.
2 Monte Carlo -body Code and Compact Object Formation
We use the Cluster Monte Carlo code (CMC; Rodriguez et al., 2022, and references therein) to study the formation of LMG BHs in GCs. CMC is a Monte Carlo -body code based on the orbit-averaged method of Hénon (1971a, b). It incorporates various relevant physics for cluster evolution, including two-body relaxation, strong three- and four-body dynamical interactions, and tidal mass loss from the galactic tidal fields. The Fewbody package is used to integrate directly strong three- and four-body gravitational encounters (Fregeau et al., 2004; Fregeau & Rasio, 2007), which includes post-Newtonian dynamics for BHs (Antognini et al., 2014; Amaro-Seoane & Chen, 2016; Rodriguez et al., 2018a, b). Single and binary star evolution is fully coupled to the stellar dynamics and is computed with COSMIC (Breivik et al., 2020), which is based on SSE (Hurley et al., 2000) and BSE (Hurley et al., 2002).
2.1 Core-collapse Supernova
BHs and NSs can form in core-collapse supernovae in CMC. The birth mass distribution of compact objects depends on the supernova mechanism and the formation of LMG BHs may be determined by effects including the explosion energy and the development timescale of instabilities (e.g., Fryer etĀ al., 2012; Belczynski etĀ al., 2012; Fryer etĀ al., 2022), or the successfulness of the supernova explosion (e.g., Kochanek, 2014). We follow the prescriptions in Fryer etĀ al. (2012) for the remnant mass and natal kick distributions of core-collapse supernovae and adopt either their ārapid modelā or ādelayed modelā in the simulations. The delayed mechanism launches an explosion in a longer timescale than the rapid mechanism and allows for a broader range of explosion energy and fallback material onto the proto-NS. This mechanism thus produces a continuous range of remnant masses across the LMG (assuming to be between and ), in contrast to the rapid mechanism which leads to a sharp transition between BHs and NSs and produces very few remnants in the mass gap (FigureĀ 4 in Fryer etĀ al. (2012) and also the ārapidā and ādelayedā histograms in FigureĀ 2 below). We adopt a maximum NS mass at following, e.g., Fryer etĀ al. (2012).222The maximum mass of NSs is still highly uncertain and depends on their equation of states (e.g. Ćzel & Freire, 2016), which allows a maximum mass between about and .
We assume that newly born NSs from core-collapse supernovae receive large natal kicks drawn from a Maxwellian distribution with a velocity dispersion (Hobbs etĀ al., 2005). The natal kicks are three-dimensional, with their directions sampled isotropically. On the other hand, the natal kicks of BHs are determined by how much material falls back post-explosion and are calculated by
where measures the fraction of the stellar envelope mass that falls back upon core collapse. Most BHs in our simulations with massive progenitor stars of have a considerable amount of fallback materials and are formed with small natal kicks (Fryer etĀ al., 2012). BHs from low-mass progenitors, however, are assumed to form with a small amount of fallback materials and receive large natal kicks similar to NSs.
2.2 Electron-capture Supernova and White Dwarf Collapse
In addition to core-collapse supernovae, NSs can form from electron-capture supernovae in our simulations (Nomoto, 1984, 1987). Electron-capture can also be triggered in massive white dwarfs (WDs) through accretion onto the WDs or mergers between two WDs (Nomoto & Kondo, 1991; Saio & Nomoto, 1985, 1998, 2004; Shen etĀ al., 2012; Schwab etĀ al., 2015, 2016; Schwab, 2021). We assume that all super-Chandrasekhar mergers where the components are Carbon-Oxygen (CO) or Oxygen-Neon (ONe) WDs collapse to NSs (e.g., Kremer etĀ al., 2023; Ye etĀ al., 2024), thus maximizing the NS formation rate through this channel. For the natal kicks of electron capture supernovae, we adopt the prescription in Kiel etĀ al. (2008) and Kiel & Hurley (2009) and sample the three-dimensional kicks from a Maxwellian distribution with a velocity dispersion . Therefore, many NSs formed in electron capture supernovae are retained in typical GCs with escape velocities of .
Here we follow the treatments of the merger-induced collapses of massive WDs from Ye etĀ al. (2024, and references therein). We summarize below the two limiting assumptions for bracketing the uncertainties in the mass loss for merger-induced collapses of super-Chandrasekhar WDs:
ā¢
Default scenario: As a lower limit, we follow the default mass prescriptions in Breivik etĀ al. (2020), where the more massive WD accretes from the donor WD during dynamical mass transfer. is the Eddington accretion limit, and is the characteristic mass-transfer timescale where is the Kelvin-Helmholtz timescale and is the dynamical timescale of the donor WD. This prescription only allows the more massive WD to accrete a small amount of mass from the donor WD, e.g., for a WD accreting for yr. The accreting WD in this scenario will collapse to an NS of gravitational mass , where is the critical mass for electron capture.
ā¢
Updated Scenario: As an upper limit, on the other hand, we assume that the mass loss is negligible during the merging and the subsequent thermal and viscous evolution phase, and the total mass of the pair of WDs is conserved up until collapse (Schwab, 2021). This upper limit allows more WD merger products to exceed the Chandrasekhar mass, thus producing a larger number of NSs and also more massive NSs. Furthermore, if the masses of the merger products exceed the maximum NS mass defined above, we assume that the remnants collapse to a BH.
2.3 Collapse of Neutron Stars
Furthermore, LMG BHs can be produced through dynamical interactions in dense star clusters. NSs engage in frequent dynamical encounters in GCs, which can lead to direct collisions with other stars or compact objects, tidal disruptions of main-sequence stars, or the formation of binaries that could subsequently merge. These collisions and mergers allow the NSs to grow in mass and collapse to BHs if exceeding the maximum NS mass. Specifically, we assume that the remnants of NSs colliding or merging with another compact object conserve the total mass of the merger components in CMC. For collisions with or tidal disruptions of main-sequence stars, we follow the treatments of NS tidal disruption events (TDEs) from Kremer etĀ al. (2022) and Ye etĀ al. (2024) to estimate the accretion rates of disruption debris of the main-sequence stars.
To briefly summarize, NS-main sequence TDEs occur when the pericenter distance between the two objects during a close encounter is smaller than the tidal disruption radius of the main-sequence star (e.g., Rees, 1988; Perets etĀ al., 2016):
where is the initial mass of the NS, and and are the initial mass and radius of the main-sequence star, respectively.
The mass accretion rate onto the NSs during TDEs is calculated as:
where , , and are the initial mass and radius of the accretion disk, and the viscous accretion time, respectively (Metzger etĀ al., 2008; Kremer etĀ al., 2022). The exponent is a free parameter determining the amount of mass transported from the outer edge of the accretion disk to the NS (e.g., Blandford & Begelman, 1999), and (e.g., Kumar etĀ al., 2008). We assume that following the results of the hydrodynamic simulations of Kremer etĀ al. (2022). This is a reasonable assumption since the relative velocity of the NSs and main-sequence stars in typical GCs is, in general, much smaller than the stellar escape velocity. We also assume that , day, and equals the NS radius (Kremer etĀ al., 2022). For more details of the TDE treatments in CMC, see Ye etĀ al. (2024).
3 Modeling Core-collapsed Globular Clusters
3.1 Typical Core-collapse Globular Clusters
We first run 13 core-collapse GC simulations up to a Hubble time (13.8Ā Gyr) with the same initial conditions using CMC to explore the formation of LMG BHs through core-collapse supernovae, merger-induced collapse of WDs, and tidal disruption of stars by NSs. These cluster simulations have initial number of stars , initial virial radius pc, metallicity , and Galactocentric distance kpc. We adopt the Kroupa standard broken power-law (Kroupa, 2001) for the initial mass function between and . The initial binary fraction is assumed to be , representative of the observed binary fraction of old GCs (e.g., Milone etĀ al., 2012).333Studies (e.g., Fregeau etĀ al., 2009) have shown that the overall hard binary fractions of GCs remain roughly constant with time. In addition, because of the large natal kicks for NSs and LMG BHs, varying the initial binary fraction will likely have negligible effects on the formation of binaries with LMG BHs. These simulations will evolve to be ātypicalā core-collapsed clusters similar to NGCĀ 6752 in the Milky Way which shows clear evidence for frequent dynamical interactions of lower-mass compact objects (e.g., single MSPs; for more details, see Ye etĀ al., 2024). Thus clusters similar to NGCĀ 6752 are ideal test grounds for studying the dynamical formation and evolution of LMG BHs. We show the surface brightness profiles and velocity dispersion profiles of all 13 simulations at Gyr in FigureĀ 1 (left panels). We focus on core-collapsed GCs in this study since they produce binaries containing NSs, low-mass BHs, or LMG BHs at higher rates than non-core-collapsed GCs (e.g., Ye etĀ al., 2020).
For these simulations, we vary the supernova prescriptions, the mass scenarios for WDāWD mergers, and the accretion efficiency of NSs tidally disrupting main-sequence stars (SectionĀ 2). These variations will affect the mass distributions of NSs and the number of LMG BHs formed. We also vary the prescriptions for the tidal capture of two WDs, and the stability criterion of binary mass transfer, . This stability criterion will influence the orbital period distribution of the compact object binaries, thereby affecting their cross sections for dynamical encounters (for more details, see Ye etĀ al., 2024). TableĀ 3.2 lists the different prescriptions adopted for each simulation.
To explore the extreme upper limit for the number of LMG BHs and their binaries formed and retained in GCs, we also run a simulation (simulationĀ 14) which truncates the standard Kroupa initial mass function at , resulting in a much smaller number of BHs with masses greater than about . Additionally, we assume that NSs born from core-collapse supernovae receive small natal kicks same as those from electron-capture supernovae. These assumptions allow for a large number of NSs and LMG BHs to be retained in the cluster. Other initial conditions of this simulation are the same as mentioned above.
3.2 Modeling NGCĀ 1851
In addition to the main grid of models which are intended to be representative of a ātypicalā core-collapsed Milky Way GC, we also run two additional higher mass simulations designed specifically to match the observed properties of NGCĀ 1851. For these models, we assume an initial number of stars and virial radius Ā pc. The metallicity of the simulations is , closely matching the observed metallicity of NGCĀ 1851 from Harris (1996, 2010 edition). We adopt a circular orbit with Galactocentric distance Ā kpc for the cluster following https://people.smp.uq.edu.au/HolgerBaumgardt/globular/, which shows that NGCĀ 1851 is on an eccentric orbit in the Milky Way with pericenter distance Ā kpc and apocenter distance Ā kpc. At distances near the pericenter distance, NGCĀ 1851 may be subject to strong tidal stripping from the Milky Way, and a simple circular orbit at the apocenter distance would not be able to capture the detailed evolution of the cluster in the Milky Wayās potential. However, simulations with Ā kpc match the observed surface brightness profile of NGCĀ 1851 at the outermost part better than those with smaller Galactocentric distances, as we will demonstrate below. This is not surprising since the cluster spends most of its orbital time near the apocenter distance, while only a short time at the pericenter distance. In addition, back scatteringāgravitational interactions with other stars scattering the potential escapers back down to a bound orbitāmay also prevent a significant fraction of objects from escaping from their host clusters even though they have enough energy to do so (e.g., Weatherford etĀ al., 2023). Simulating the detailed evolution of the clusterās tidal boundary is beyond the scope of this paper, and the location of the tidal boundary likely does not have a significant effect on the LMG BHs, which mostly concentrate at the cluster cores due to mass segregation. Other initial cluster properties (e.g., binary fraction, initial mass function, stellar distribution profile) are the same as the simulations in the main grid (simulations 1-13 above). For these two simulations, we again adopt different prescriptions that will affect the mass of the NSs and BHs for comparison (simulationsĀ A and B in TableĀ 3.2). Furthermore, given the observed eccentric orbit of NGCĀ 1851, we include simulationĀ C with Galactocentric distance Ā kpc as a limiting case. All the other initial conditions of this simulation are the same as in simulationĀ B.
Model | SN | TDE | WDTC | qcrit | |||||||
1 | 1.97 | 5 | 968 | 18 | 1 | 0 | rapid | default | - | - | default |
1.90 | 2 | 905 | 17 | 2 | 0 | delayed | default | - | - | default | |
1.99 | 7 | 917 | 21 | 3 | 1 | rapid | updated | - | - | default | |
1.89 | 5 | 862 | 16 | 0 | 0 | delayed | updated | - | - | default | |
5 | 2.01 | 3 | 889 | 17 | 2 | 0 | rapid | default | 0.2 | ā | default |
2.02 | 3 | 970 | 20 | 1 | 0 | delayed | default | 0.2 | ā | default | |
1.97 | 12 | 904 | 12 | 0 | 0 | rapid | updated | 0.2 | ā | default | |
2.04 | 9 | 937 | 21 | 2 | 1 | delayed | updated | 0.2 | ā | default | |
9 | 2.07 | 4 | 961 | 22 | 1 | 0 | rapid | default | 0.2 | ā | updated |
1.96 | 1 | 1077 | 13 | 1 | 0 | delayed | default | 0.2 | ā | updated | |
2.01 | 15 | 955 | 21 | 2 | 1 | rapid | updated | 0.2 | ā | updated | |
1.98 | 10 | 968 | 12 | 2 | 0 | delayed | updated | 0.2 | ā | updated | |
2.05 | 10 | 905 | 16 | 1 | 1 | delayed | updated | 0.0 | ā | updated | |
2.03 | 13 | 3181 | 10 | 6 | 1 | delayed | updated | 0.2 | ā | updated | |
A | 3.79 | 17 | 816 | 20 | 0 | 0 | rapid | default | - | - | default |
3.80 | 24 | 970 | 22 | 1 | 1 | delayed | updated | 0.2 | ā | updated | |
2.23 | 15 | 936 | 20 | 2 | 1 | delayed | updated | 0.2 | ā |
Note. ā From left to right, the columns are the total mass of the cluster at present, the number of BHs retained at present, the number of NSs retained at present, and the number of compact object binaries (NSāWD, NSāNS, and BHāNS) at present. The present-day values for simulations 1-14 are averages between 11 and 13.8Ā Gyr in each cluster simulation, and between 10 and 13Ā Gyr for simulationsĀ A, B, and C.
āThe last five columns indicate the various physical processes included in the simulations and relevant to forming LMG BHs. āSNā shows the supernova model used. āā shows the mass scenario assumed for the merger-induced collapse of two WDs (SectionĀ 2). āTDEā indicates whether the simulations include the tidal disruption of main-sequence stars by NSs and the subsequent accretion of the disrupted debris by the NSs (see also SectionĀ 2.3). The numbers show the -parameter values assumed for EquationĀ 3. corresponds to the highest accretion efficiency. āWDTCā indicates if WDāWD tidal capture is turned on in the simulations. The inclusion of WDāWD tidal capture can slightly increase the WD merger rates. āqcritā shows the critical mass prescription used for stable mass transfer (default: Hurley etĀ al., 2000 and Hjellming & Webbink, 1987; updated: Belczynski etĀ al., 2008). The updated prescription for āqcritā allows for more stable mass transfer, thus generally producing wider binaries that are easier to disrupt by gravitational encounters. For more details, see Ye etĀ al. (2024).
āNote that simulationĀ 14 is an extremely optimistic model where most NSs and LMG BHs receive small natal kicks and are retained in the cluster. The initial mass function of this model is also truncated at to prevent the formation of massive BHs.
āSimulationĀ C has the same initial conditions as simulationĀ B except for the Galactocentric distance . SimulationĀ B has Ā kpc while simulationĀ C has Ā kpc.
FigureĀ 1 (right panels) shows comparisons between the model and observed surface brightness profiles and velocity dispersion profiles of NGCĀ 1851. We exclude very luminous stars with luminosity to avoid large fluctuations in the model surface brightness profiles, and we calculate the model velocity dispersion profiles with only the giant stars and upper main-sequence stars with (the main-sequence turnoff mass of the simulations is ). The orange profiles correspond to simulation A in TableĀ 3.2 (which assumes the rapid supernova prescription and the default mass scenario for WDāWD mergers). The green profiles are for simulation B in the same table (which assumes the delayed supernova prescription and the updated WD mass scenario). Both simulations produce similar surface brightness and velocity dispersion profiles at the present day because of the same initial conditions used. We note that the model surface brightness profiles do not overlap exactly with the observed ones. However, our goal is not to produce a best-fit NGCĀ 1851 model in this study, but to explore the formation pathways of binaries similar to the NGCĀ 1851E binary in models that match the observed properties of NGCĀ 1851 reasonably closely.
TableĀ 3.2 summarizes the present-day properties of the 14+2 GC simulations and their prescriptions for forming LMG BHs. The present-day values are averaged between 11 and 13.8Ā Gyr for simulations 1-14, while between 10 and 13Ā Gyr for simulations A-B (the observed age of NGCĀ 1851 is Ā Gyr; VandenBerg etĀ al., 2013). All clusters have reached core-collapsed at late times, which is shown by the small number of BHs retained (e.g., Kremer etĀ al., 2020a, and references therein). In the following sections, we will explore the contributions from various pathways to the formation of binaries and transient sources containing LMG BHs.
4 Formation of Lower-mass-gap Objects
In this section, we quantify the effects of the rapid and the delayed models of core-collapse supernovae from Fryer etĀ al. (2012) (SectionĀ 2.1), accretion- and merger-induced collapses of WDs (SectionĀ 2.2), as well as the collapses of NSs through mergers and TDEs (SectionĀ 2.3), on the population of LMG BHs in GCs.
FigureĀ 2 shows the mass distributions of BHs and NSs formed over a Hubble time in simulations 1-12. The simulations are divided into four groups that utilize the rapid and delayed supernova prescriptions (Fryer etĀ al., 2012) as well as the default and updated mass scenarios for the mergers between a pair of WDs (TableĀ 3.2). Simulations that adopt the delayed supernova prescription produce a significant number of objects in the LMG ( per cluster over a Hubble time), in contrast to the ones using the rapid prescription ( per cluster over a Hubble time).
Despite the large number of LMG BHs born, most of them are not retained in the host clusters. In our simulations (except the extremely optimistic simulationĀ 14), NSs up to receive large natal kicks of hundreds of , and BHs smaller than about receive fallback kicks at birth similar to the NS kicks (SectionĀ 2). The upper panel of FigureĀ 3 shows the natal kick distributions for massive NSs and low-mass BHs in simulations 1-12. More than of the compact objects between and have natal kicks . Since the escape velocities of typical GCs are generally , most of the LMG BHs formed from the delayed prescriptions are ejected from their host clusters at birth. This leads to the retention of only a few LMG BHs from the delayed supernovae. The bottom panel of FigureĀ 3 shows the mass distributions of NSs and BHs of the same 12 simulations at the present day. The similarity of the remnant mass distributions between the simulations using the rapid and delayed prescriptions demonstrates the negligible effects of the supernova prescriptions on the LMG populations in typical GCs, if LMG remnants have large natal kicks. TableĀ 4 also shows the number of LMG BHs formed in supernovae that are retained in the clusters ( per cluster for typical Milky Way core-collapsed clusters; simulation 1-13). If all NSs and LMG BHs receive small natal kicks, however, hundreds of LMG BHs can be retained in a cluster (simulationĀ 14). For more massive clusters similar to NGCĀ 1851 (simulations A and B), the number of retained LMG BHs is larger due to the clusterās higher escape velocity.
Columns 2-5 in TableĀ 4 show the contributions to the formation of LMG BHs from various channels. These numbers only include the retained remnants and the mergers in the clusters. Overall, mergers between NSs and WDs produce the largest number of LMG BHs compared to other channels (except when TDE accretion is almost conservative and when all NSs and LMG BHs receive low natal kicks), with the numbers comparable to or much larger than the combined total from other channels. Those involving two NSs contribute to of the numbers formed by the NSāWD merger channel in most cases, with variations for different simulations. This is not surprising since the number of massive WDs in GCs is significantly larger than that of NSs from stellar evolution, even though NSs are more massive.
For simulations 1-13 that adopt the delayed supernova prescription, LMG BHs are retained. This number can be comparable to that from the NSāNS merger channel in simulations with the default mass scenario for WD mergers, but it is much smaller than either the NSāWD or NSāNS channel when the updated mass scenario is adopted. On the other hand, for more massive clusters (e.g., clusters like NGCĀ 1851; simulation B) the number of retained LMG BHs formed through delayed supernovae can be several times higher due to the larger cluster escape velocity, resulting in a contribution similar to that of the NSāWD merger channel. We can expect that in even more massive clusters or with smaller natal kicks, more LMG BHs could be produced by supernovae rather than forming through the collapse of NSs from dynamical interactions. We will explore different supernova prescriptions and compact object natal kicks in future studies.
NS-main-sequence star TDEs generally have minor contributions to forming LMG BHs even with relatively high accretion efficiency of the tidally disrupted debris (TableĀ 4). Simulations 5-12 assume that in EquationĀ 3, corresponding to accreted onto the NS in Ā seconds from the disruption of a main-sequence star (Kremer etĀ al., 2022). This amount of mass is insufficient to cause NSs with typical masses of about to collapse to BHs. Only with the largest accretion rate when (corresponding to accreted in Ā seconds from the disruption of a main-sequence star) can many NSs collapse to low-mass BHs from the tidal disruption channel. This is demonstrated in TableĀ 4 by simulation 13, where the TDE channel becomes the dominant channel for producing LMG BHs.
Note. ā Columns 2-5 show the delayed supernova formation and the three dynamical formation pathways of BHs in the LMG (). From left to right, it is the number of LMG BHs from the delayed supernova prescription, number of NS collapsing to BHs through NS-main-sequence star TDEs, binary NS mergers, and NSāWD mergers, respectively.
āColumns 6-11: the number of various mergers (binary BHs, binary NSs, BHāNS, BHāWD, NSāWD, and binary WDs) in each simulation over a Hubble time. Columns 12-13 show the fraction of BHāNS and binary BH mergers that contain at least one LMG BHs.
āColumns 14-16 are the collisions and tidal disruptions between WDs and main-sequence stars, NSs and main-sequence stars, as well as BHs and main-sequence stars, respectively. The last column shows the fraction of BH-main-sequence star collisions/TDEs that involve a BH in the LMG.
āSimulationĀ 14 shows an extreme case where all NSs and LMG BHs receive small natal kicks the same as those from electron-capture supernovae, and the initial mass function of the simulation is truncated at .
āSimulationĀ C shows a limiting case with Galactocentric distance Ā kpc. Other initial conditions are the same as in simulationĀ B.
5 Lower-mass-gap Binaries
Having established that LMG BHs can form in GCs through various mechanisms, we now explore the subsequent dynamical evolution of these objects, in particular the dynamical formation of LMG binaries. In general, only a very small number of NSāNS and BHāNS binaries are present at late times for all cluster simulations, as shown in TableĀ 3.2 (average of 6 and up to about 10 in the extremely optimistic simulationĀ 14). This is roughly consistent with the fact that, despite the hundreds of MSPs observed in Milky Way GCs (and found in simulations, e.g., Ye etĀ al., 2020, 2022), only two pulsar-NS (or low-mass BH) binaries have been definitively identified. Both are in massive, core-collapsed, or near core-collapsed clusters (M15 and NGC 1851). We list the properties of the MSP NGCĀ 1851E and its binary in TableĀ 3, together with those of the confirmed or candidate NSāNS binaries in all Milky Way GCs for reference.
However, different mass prescriptions for compact objects can significantly alter the present-day mass distribution of NSāNS and BHāNS binaries, as demonstrated in FigureĀ 4 for simulations 1-12. The delayed supernova prescription is more likely to produce BHāNS binaries with BH mass compared to the rapid prescription. Still, both supernova prescriptions yield similar numbers of binaries with massive NSs and LMG BHs (FigureĀ 4, upper left panel). The most significant change in the BH/NSāNS binary mass distribution arises from the updated mass scenario for WD merger and the subsequent collapse of the heavy remnant, conserving mass in both evolutionary stages. As discussed above, this updated mass scenario leads to more massive NSs and more LMG BHs, which is illustrated by the broader span of secondary masses and the increased occurrence of binaries with massive NSs and LMG BHs in the right panels of FigureĀ 4. For simulations 1-12, approximately of the primaries in all BHāNS binaries containing an LMG BH () between 11 and 14Ā Gyr form from NSāWD mergers, about from NSāNS mergers, and around from NSāMS TDEs. If we also include binaries containing massive NS primaries in the range of , then approximately of the massive NSs and LMG BHs form from NSāWD mergers, around from NSāNS mergers, about from NSāMS TDEs, and roughly from WDāWD mergers (not directly forming LMG BHs but massive NSs). Overall, NSāWD mergers contribute the most to the production of massive NS and LMG BH in binaries, followed by NSāNS mergers, in typical core-collapsed GCs similar to NGCĀ 6752. FigureĀ 6 illustrates the dynamical formation of a NSāNS binary from simulationĀ 12 where the primary NS forms from a NSāWD collision.
Cluster | PSR | P | References | ||||||||
[] | [arcmin] | [arcmin] | [ms] | [day] | [] | [] | |||||
NGC 1851 | 2.8 | 0.09 | 0.51 | 15 | NGC 1851E | 5.596 | 7.448 | 0.708 | (1) | ||
NGC 7078 (M15) | 5.2 | 0.14 | 1.00 | 13 | M15C | 30.529 | 0.335 | 0.681 | (2) | ||
āNGC 1851 | 2.8 | 0.09 | 0.51 | 15 | NGC 1851A | 4.991 | 18.785 | 0.888 | (3) | ||
āNGC 6544 | 0.8 | 0.05 | 1.21 | 3 | NGC 6544B | 4.186 | 9.957 | 0.747 | 1.3655 | 1.2064 | (4) |
āNGC 7099 (M30) | 1.2 | 0.06 | 1.03 | 2 | M30B | 12.990 | 6.216 | 0.879 | (5) | ||
āNGC 6838 (M71) | 0.38 | 0.63 | 1.67 | 5 | M71D | 100.670 | 10.939 | 0.640 | - | 1.63 | (6) |
Note. ā From left to right, the columns show the name of the cluster, cluster mass from https://people.smp.uq.edu.au/HolgerBaumgardt/globular/, the observed cluster core and half-light radii from Harris (1996, 2010 edition), the observed number of pulsars from https://www3.mpifr-bonn.mpg.de/staff/pfreire/GCpsr.html, the name of the pulsarāNS/BH binary or candidate, the pulsarās spin period, binary orbital period, binary eccentricity, mass of the pulsar, mass of the companion, and the reference for the binary properties, respectively. (1) Barr etĀ al. (2024); (2) Jacoby etĀ al. (2006); (3) Freire etĀ al. (2004), Freire etĀ al. (2007), and Ridolfi etĀ al. (2019) (4) Lynch etĀ al. (2012); (5) Ransom etĀ al. (2004) and Balakrishnan etĀ al. (2023); (6) Pan etĀ al. (2021), https://www3.mpifr-bonn.mpg.de/staff/pfreire/GCpsr.html, and https://fast.bao.ac.cn/cms/article/65/.
āAll of the cluster listed are core-collapsed or nearly core-collapsed (except M71). Note that even though NGCĀ 1851 is shown as non-core-collapsed in Harris (1996, 2010 edition), it has a very small core (e.g., Miocchi etĀ al., 2013 and Harris, 1996, 2010 edition) and a very high central density (Baumgardt & Hilker, 2018), similar to core-collapsed clusters.
āA indicates a binary candidate.
Similarly, both NGCĀ 1851-like simulations produce comparable numbers of NSāNS and BHāNS binaries at present (TableĀ 3.2). Simulation A has zero, while simulation B has about 2 NS/BHāNS per snapshot, roughly consistent with the current pulsar timing observation. Note that NGCĀ 1851E-like pulsar binaries with orbital periods longer than about a week have selection biases similar to other pulsar binaries with massive companions (around ) and can be detected at around apoastron. Therefore, if a large number of them exist in GCs, they would potentially be detectable.444However, there are strong selection biases against finding NSāNS binaries with orbital periods shorter than about two days (similar to the M15C binary) due to the effects of large orbital accelerations on pulsar searches. It would be difficult to determine the underlying population of M15C-like binaries in GCs.
On the other hand, the component mass distribution of the NSāNS and BHāNS binaries differs between the two simulations with different mass scenarios for WD mergers and supernova prescriptions. FigureĀ 5 shows the primary and secondary masses of all NSāNS and NSāBH binaries between 10 and 13Ā Gyr of the clustersā evolution from both NGC 1851-like models. The component masses of the NGCĀ 1851E binary are plotted for comparison. Simulation A produces no BHāNS binaries with primary mass in the LMG. In contrast, for () of the time steps between and Ā Gyr, the NSāNS or BHāNS binaries in simulation B have primary mass in the range (), including binaries with masses similar to the NGCĀ 1851E binary (bottom panel of FigureĀ 5). Furthermore, simulation B also produces more massive NSs that can better match the mass of NGCĀ 1851E. This figure suggests that the delayed supernova prescription (or, more generally, the formation of LMG BHs from massive star evolution and their retention in dense star clusters) is important for producing LMG binaries in dense and massive clusters similar to NGCĀ 1851. This is because massive clusters can retain more LMG BHs due to their larger escape velocities (TableĀ 4). Furthermore, it also indicates that conserving mass during WDāWD mergers (the updated mass scenario in SectionĀ 2) is a key factor for clusters to produce NGC 1851E-like binaries containing an LMG BH and a massive NS companion (see also, e.g., Ye etĀ al., 2024, their FigureĀ 9).
Many NS/BHāNS binaries between 10 and 13Ā Gyr of simulation B have component masses and orbital properties similar to the NGCĀ 1851E binary. In particular, for binaries with primary masses in the range of and secondary masses in the range of (approximately the same as the components in the NGCĀ 1851E binary), the mean and median eccentricities are about 0.7, and the mean and median orbital periods are about 31 and 13 days, respectively 555Note that these values come from only one simulation. It is possible that the values could change slightly with more simulations using the same initial conditions but different random seeds..
Among the NS/BHāNS binaries in simulation B between and Ā Gyr whose primary mass is in the range of (), () of the time the secondary NS is an MSP (including MSPs formed through NS-main-sequence star TDEs; Kremer etĀ al., 2022; Ye etĀ al., 2024). We assume that NSs produced in WDāWD mergers are born as young pulsars (Kremer etĀ al., 2023) rather than MSPs (Ye etĀ al., 2024). These young pulsars are expected to spin down quickly in Ā Myr, and are likely undetectable in the present day.
Overall, our NGCĀ 1851-like models can produce NSāNS or BHāNS binaries similar to the NGCĀ 1851E binary. A primary in the LMG most likely forms in a delayed supernova (approximately of all binaries containing an LMG primary between 10 and 13Ā Gyr) or through the merger between an NS and a massive WD (around ). Only about of these LMG primaries form from NSāNS mergers. By also including binaries containing massive NS primaries in the range of , the fraction from the delayed supernovae changes to approximately , the fraction from NSāMS TDEs increase to about , and WDāWD mergers contribute around , with other fractions remaining mostly the same. Thus, the LMG or massive NS companion of NGCĀ 1851E is less likely to be formed in NSāNS mergers. These fractions are very different from those mentioned above for simulations 1-12 due to the difference in the cluster mass and escape velocity. FigureĀ 6 shows an example of the formation of one of the LMG BHāNS binaries in simulation B.
6 Gravitational Wave Sources and Luminous Transients with Lower-mass-gap Black Holes
6.1 Gravitational Wave Sources
Some LMG BHs can be paired up with another compact object through dynamical encounters in the clusters and inspiral to produce GW signals in subsequent mergers observable by LVK.
FigureĀ 7 shows the component masses of BHs and NSs in all merging binaries over a Hubble time from simulations 1-12. Three GW sources GW230529 (The LIGO Scientific Collaboration etĀ al., 2024), GW200115 (Abbott etĀ al., 2021a), and GW190814 (Abbott etĀ al., 2020) are shown as orange crosses and error bars for comparison. Overall, both supernova prescriptions lead to similar numbers of binary BH mergers (), while the delayed supernova prescription produces slightly more (a factor of ) BHāNS and NSāNS mergers (upper panel). This small increase is consistent with most LMG BHs from the delayed supernovae being ejected from the clusters by their natal kicks (SectionĀ 2.1). Simulations with the updated mass scenario for WD mergers form times more BHāNS and NSāNS mergers compared to those with the default mass scenario (when all other initial conditions and supernova prescriptions are the same) as shown in the lower panel of the same figure. This increase results from the larger number of massive NSs and WDs formed because of the conservation of mass in WD mergers (also see SectionĀ 2.3). Therefore, simulations with both the delayed supernova prescription and the updated WD mass scenario produce the largest number of NSāNS ( per core-collapsed cluster up from ) and BHāNS ( per core-collapsed cluster up from ) mergers. TableĀ 4 shows the number of various compact object mergers over a Hubble time and the fraction containing LMG BHs in each simulation.
In general, only a small fraction of binary BH mergers from GCs (a few percent) include LMG BHs between and . However, of merging BHāBH from a typical core-collapsed GC can contain at least one LMG BH if NSs are able to accrete very efficiently from TDEs of main-sequence stars and collapse to BHs (simulation 13 in TableĀ 4). These BHs would be fast-spinning and may be detectable by LVK. On the contrary, of BHāNS mergers involve a BH in the LMG for most simulations in TableĀ 4. The recent detection of GW230529, a merger between a LMG BH and an NS from the fourth observing run of LVK, suggests that about of BHāNS mergers may include a component in the LMG (The LIGO Scientific Collaboration etĀ al., 2024). This fraction is similar to what we find for BHāNS mergers in core-collapsed clusters, but the merger rates of GW230529-like systems, (The LIGO Scientific Collaboration etĀ al., 2024), are much larger than the rates expected from GCs. Assuming that about of GCs are core-collapsed (similar to the percentage in the Milky Way) and the GC number density is about (Rodriguez etĀ al., 2016a), the merger rates of all BHāNS mergers from GCs are roughly , even when adopting the delayed supernova prescription and the updated WD mass scenario (the actual rate is probably even lower; Ye etĀ al., 2020, see). The extreme upper limit of the BHāNS merger rates is a factor of a few higher and roughly comparable to the BHāBH merger rates, assuming that all NSs and BHs receive small natal kicks, and that not many BHs with masses greater than are formed in clusters (simulationĀ 14). Therefore, even though we can roughly reproduce the observed MSPāLMG BH in NGCĀ 1851, the estimated merger rate of BHāNS is still likely orders of magnitude lower than inferred from the most recent LVK data release (; Abbott etĀ al., 2023).
6.2 Transients from Stellar Collisions and Tidal Disruptions
NSs and stellar-mass BHs also dynamically interact with the luminous stars in GCs, as the compact object populations are mixed with the numerous upper main-sequence stars and giant stars in the cluster cores (e.g., Kremer etĀ al., 2020b; Ye etĀ al., 2022). This is directly evidenced by the growing number of observed redback MSPs (https://www3.mpifr-bonn.mpg.de/staff/pfreire/GCpsr.html) and BH binaries with luminous companion stars (e.g., Giesers etĀ al., 2019) in GCs. Occasionally, the luminous stars pass sufficiently close to the NSs or BHs to be tidally disrupted or directly collide with them.
We record the number of main-sequence star collisions or TDEs with NSs and BHs over a Hubble time in the simulations in TableĀ 4. In typical core-collapsed clusters with mass of about (simulations 1-13), the number of NSāmain-sequence TDEs is comparable to the number of BHāmain-sequence TDEs, different from the non-core-collapsed clusters where BH TDEs are two orders of magnitude more frequent (e.g., Kremer etĀ al., 2019). The more massive, NGCĀ 1851-like simulationsĀ A andĀ B have a smaller number of NS TDEs compared to their BH TDEs, by a factor of about 2ā7 depending on the supernova prescription and WD mass scenario. Overall, simulations using the rapid supernova prescription and the default mass scenario for WD merger produce the smallest numbers of TDEs. Among the BHāmain-sequence TDEs from simulations 1-12, about involve a BH in the LMG, with the largest fraction coming from models that adopt the delayed supernova prescription and/or the updated mass scenario. The fractions involving a LMG BH are similar for the NGCĀ 1851-like simulationsĀ A andĀ B. Furthermore, at most of BHāmain-sequence TDEs can be carried out by LMG BHs in typical core-collapsed clusters if we assume that all mass bound to the NSs during NSāmain-sequence TDEs are accreted by the NS, which results in more NSs collapsing to LMG BHs (simulation 13 in TableĀ 4). These TDEs by LMG BHs may give rise to bright X-ray and gamma-ray flares which could resemble ultra-long gamma-ray bursts (Perets etĀ al., 2016). They may also produce optical signals of luminosity that could be detectable by optical transients surveys such as the Zwicky Transient Facility (e.g., Kremer etĀ al., 2019).
7 Discussion
One of the main uncertainties of the formation and evolution of LMG BHs in dense star clusters lies in the number of systems retained, which is determined by the supernova mechanism and the kicks low-mass BHs and NSs receive at birth. The escape velocity of typical GCs () imposes an upper limit on the retention fraction, depending on the natal kick distribution. For NS natal kicks estimated from isolated pulsar observations in the Galactic field (e.g., Hobbs etĀ al., 2005, adopted in this study), few massive NSs and low-mass BHs (less than about ) are retained (TableĀ 4), even though are formed in a typical GC assuming the delayed supernova mechanism (Fryer etĀ al., 2012). More recent studies, either based on observed NS binaries (e.g., OāDoherty etĀ al., 2023) or on three-dimensional supernova simulations and semi-analytical models (e.g., Mandel & Müller, 2020), have instead suggested that NSs may receive smaller natal kicks than previously estimated (e.g., FigureĀ 1 in OāDoherty etĀ al., 2023, and references therein). If this is the case, up to times more NSs may be retained in GCs, considering also that most massive stars are born in binaries (Sana etĀ al., 2012; Moe & Di Stefano, 2017). Assuming that LMG BHs, if formed, also receive small natal kicks, an order of magnitude more LMG BHs could be retained in GCs. Together with the larger number of retained NSs, this might lead to a factor of a few more binaries containing LMG BHs in the clusters in the present day. Future (non)observations of LMG binaries will shed more light on the supernova and dynamical formation of these objects in dense star clusters.
We showed that for typical core-collapsed GCs with masses of about at present (e.g., NGCĀ 6752), LMG BHs more likely form through dynamical encounters, especially mergers between NSs and WDs. On the other hand, for more massive core-collapsed clusters with present-day masses of about (e.g., NGCĀ 1851), core-collapse supernovae through stellar evolution and dynamical interactions (e.g., NSāWD mergers) likely contribute similar numbers of LMG BHs (see also SectionĀ 4 and TableĀ 4). Additionally, if NSs can accrete nearly all the disrupted stellar debris during TDEs of main-sequence stars, this channel could contribute the most to the formation of LMG BHs. The above suggests that massive, non-core-collapsed clusters may also retain a small number of LMG BHs from massive star evolution.
Another main uncertainty lies in the mass accretion efficiency and mass loss rate of WDāWD mergers. Although zero LMG BHs in the simulations formed directly from the merger-induced collapse of two WDs, it has significant impacts on producing massive NSs, as shown in SectionĀ 5. For example, many secondary NSs larger than shown in FigureĀ 5 originate from WDāWD mergers, and they are important for matching the mass of NGCĀ 1851E. Significant mass loss during the mergers of two WDs would instead indicate that more massive NSs are formed and retained in the clusters compared to those formed in electron-capture supernovae. This may be due to smaller natal kicks in core-collapse supernovae, as discussed earlier, or because the NSs accrete more mass during previous mass transfer (e.g., through NSāMS TDEs).
The LMG BHs, regardless of their formation channel, could subsequently pair with an MSP or another NS/BH through dynamical encounters, most likely in massive, dense, core-collapsed clusters like M15 and NGCĀ 1851. Core-collapsed clusters probably have only a few BHs remaining (e.g., Kremer etĀ al., 2018), allowing NSs to mass segregate to the cluster cores and frequently engage in dynamical interactions (Ye etĀ al., 2019, their FigureĀ 5). Therefore, the formation of LMG BHs through the collapse of NSs and the formation of NSāNS and NSāBH binaries (Ye etĀ al., 2020) are boosted in these dense core-collapsed clusters, and it is not surprising that almost all detected MSPāNS/BH or their candidates are found in (near)core-collapsed clusters (TableĀ 3).
Similarly, the formation of MSPs (Ye etĀ al., 2019, 2024) and young radio pulsars (Kremer etĀ al., 2023) is also greatly enhanced in the dense, core-collapsed environment. The large observed number of MSPs (330 pulsars in 44 Milky Way clusters observed to date; https://www3.mpifr-bonn.mpg.de/staff/pfreire/GCpsr.html), although seemingly contradictory, is still consistent with the small number of confirmed NSāNS/LMG BHs in Milky Way GCs (M15C and NGCĀ 1851E). This is because the formation of NSāNS or NSāBH binaries is limited by the dynamical encounter rates of two NSs (single or in binary), which roughly scale with the square of the NS number density in dense star clusters. On the other hand, the formation rate of MSP binaries in dense star clusters scales with the number density of the NSs and the number density of either WDs or luminous stars, which are orders of magnitude higher than the NS number density. For example, the encounter rate between a single and a binary star can be written as
where is the number densities of the target binaries and is the total number of singles of interest. is the encounter cross section, and is the velocity dispersion. Thus orders of magnitude more observed MSPs in GCs than NSāNS/BH binaries is expected. Likewise, the formation rate of merging NSāNS and BHāNS binaries is also much lower than that of MSPs and the inferred rate from GW detections (see SectionĀ 6.1 above; also Ye etĀ al., 2020; Abbott etĀ al., 2023; The LIGO Scientific Collaboration etĀ al., 2024).
8 Conclusions
In this study, we have explored for the first time the formation and evolution of LMG BHs (in the mass range around ) and their binaries through the interplay between massive star evolution and dynamical interactions in dense star clusters. Motivated by the recent detection of a pulsar binary with a potential LMG companion in the massive and dense Milky Way cluster NGCĀ 1851 (Barr etĀ al., 2024), we model this cluster together with typical core-collapsed GCs using our Monte Carlo code CMC, which self-consistently takes into account stellar (and binary star) evolution as well as all -body dynamics.
Our simulations demonstrate the effects of two main physical mechanisms on the formation of LMG BHs. The first is the supernova core collapse mechanism (i.e., rapid vs delayed collapse; Fryer etĀ al., 2012). The second is dynamical interactions resulting in the collapse of massive compact merger remnants into low-mass BHs (including NSāWD mergers, NSāNS mergers, and tidal disruption of main-sequence stars by NSs). We consider two possible outcomes for mergers of massive WDs (our updated scenario conserves mass, while the default in CMC does not) since WDāWD mergers may produce massive NSs that can later collapse to low-mass BHs in subsequent dynamical interactions.
The delayed supernova prescription produces LMG BHs in a typical core-collapsed GC (while the rapid prescription produces none). However, with the large natal kicks they likely receive, only a few LMG BHs are retained in the clusters after formation. More massive GCs similar to NGCĀ 1851 can retain a few tens of these LMG BHs born in delayed supernovae because of the clusterās deeper potential well and the larger number of compact remnants formed initially.
Most of the LMG BHs that form in our models through the collapse of massive NSs originate from NSāWD mergers. In simulations adopting the delayed supernova prescription or the updated WD mass scenario, NSāWD mergers contribute more than twice as many LMG BHs as NSāNS mergers. This is expected since there are many more massive WDs in GCs than NSs and even though most WDs have smaller masses. Only in simulations that adopted the rapid supernova prescription and the default WD mass scenario are there comparable numbers of LMG BHs from NSāWD and NSāNS mergers. In addition, at most a few LMG BHs are produced by NSs tidally disrupting main-sequence stars and accreting the debris. However, if NSs can accrete very efficiently during TDEs (i.e., almost all the mass of the disrupted main-sequence star), then this channel could dominate the production of LMG BHs in core-collapsed clusters (e.g., Kremer etĀ al., 2022).
Among all formation channels, NSāWD mergers can produce of LMG BHs (with an upper limit of ) in typical core-collapsed GCs given the supernova (Fryer etĀ al., 2012) and natal kick prescriptions (Hobbs etĀ al., 2005) adopted in our models. On the other hand, the number of retained LMG BHs from delayed supernovae are comparable to those from NSāWD mergers in more massive core-collapsed clusters similar to NGCĀ 1851. This suggests that LMG BHs formed in supernovae may be significant in more massive clusters, or if they (and NSs) receive smaller natal kicks as indicated by recent studies (e.g., Mandel & Müller, 2020; OāDoherty etĀ al., 2023) compared to the previous estimates from isolated pulsar observations (e.g., Hobbs etĀ al., 2005). We will explore different supernova and natal kick prescriptions more systematically in future studies.
Our NGCĀ 1851-like models typically have a few BHāNS and NSāNS binaries at present, roughly consistent with the one confirmed system in NGCĀ 1851 (TableĀ 3 and Barr etĀ al., 2024) considering that binaries similar to the NGCĀ 1851E binary with orbital periods longer than about a week are not strongly affected by selection biases during pulsar searches. Future (non)detection of BHāNS and NS-NS binaries in GCs would help better constrain supernova models and natal kicks of NSs and LMG BHs. Among the late-time snapshots between 10 and 13Ā Gyr of the clusterās evolution, about 50% (65%) contain NSāNS/BH binaries with primary masses in the range (), and about 30% (45%) of the time these binaries have an MSP companion. Some of the NSāNS or NSāBH binaries in our models have both component masses similar to those inferred for NGCĀ 1851E. The potential LMG companion of NGCĀ 1851E most likely formed in a core-collapse supernova or a NSāWD merger. Simulations that assume our updated mass scenario, where the total mass is conserved for WDāWD mergers (e.g., Schwab, 2021), better reproduce the mass of NGCĀ 1851E ().